OpenStudy (darkigloo):

if y=10^(x^2-1), what is y' ?

zepdrix (zepdrix):

Have you learned the shortcut rule for differentiating something of the form,\[\Large\bf\sf a^x\]Or no?

OpenStudy (shamil98):

http://www.mathsisfun.com/calculus/derivatives-rules.html enjoy.

OpenStudy (darkigloo):

ohh..i see a^x(ln a). guess i'll plug those in.

zepdrix (zepdrix):

Yah that'll get you on the right track :) Then don't forget to apply the chain rule after that!!

OpenStudy (darkigloo):

i got the answer. thanks.

OpenStudy (darkigloo):

(2x) (ln10) 10^(x^2-1)

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

\[\ln y=\left( x^2-1 \right)\ln 10\] \[\frac{ y' }{y }=2x \ln 10,y'=y*2x \ln 10=10^{x^2-1}2x \ln 10\]

OpenStudy (darkigloo):

Thank you all. :)

OpenStudy (anonymous):

yw